Ashley is making punch for a party. The recipe requires twice as much orange juice as cranberry juice and 8 times as much ginger ale as cranberry juice. If she plans to make 176 onces of punch, how much of each ingredient should she use.
Answer by nerdybill(7384)

Ashley is making punch for a party. The recipe requires twice as much orange juice as cranberry juice and 8 times as much ginger ale as cranberry juice. If she plans to make 176 onces of punch, how much of each ingredient should she use.
.
Let x = onces of cranberry juice
then
2x = onces of orange juice
8x = onces of ginger ale
.
x + 2x + 8x = 176
11x = 176
x = 176/11
x = 16 onces (cranberry juice)
.
Orange juice:
2x = 2(16) = 32 ounces
.
Ginger ale:
8x = 8(16) = 128 ounces